What is the molarity of a 85% H2SO4 solution with a density of 1.2 g/mol?


1
Expert's answer
2021-02-03T07:23:05-0500

Density = 1.2 g/mL


Solution:

85% mass of H2SO4 means 85 g of H2SO4 in 100 g solution.


Density of solution = Mass of solution / Volume of solution

Hence,

Volume of solution = Mass / Density = (100 g) / (1.2 g mL-1) = 83.33 mL = 0.08333 L


Molarity = Moles of solute / Volume of solution = Mass / (Molar mass × Volume of solution)

The molar mass of H2SO4 is 98 g mol-1.

Molarity = (85 g) / (98 g mol-1 × 0.08333 L ) = 10.4086 mol/L = 10.4 M

Molarity of a 85% H2SO4 solution is 10.4 M.


Answer: The molarity of a 85% H2SO4 solution is 10.4 M.


OR:


Molarity = (w × d × 10) / M

where,

w = amount of solute

d = density of solution

M - molecular mass of solute


Molarity = (85 × 1.2 × 10) / 98 = 10.408 M

Molarity of a 85% H2SO4 solution is 10.4 M.


Answer: The molarity of a 85% H2SO4 solution is 10.4 M.